A new cruise ship line has just launched 3 new ships. The Pacific, Caribbean, and Mediterranean. The Caribbean has 32 more deluxe rooms than the Pacific.The Mediterranean has 37 fewer deluxe rooms than twice the number of deluxe rooms on the Pacific. Find the number of deluxe rooms for each of the ships if the total number of deluxe rooms for the 3 ships is 607.
1
Expert's answer
2012-07-17T08:09:08-0400
Let Pacific has x rooms. Then Caribean has x+32 and Mediterranean has 2x-37.
So& x+x+32+2x-37=607
4x=607+37-32=612
x=153 - Pacific
153+32=185 - Caribean
153*2-37=269 -Mediterranean